What is the latitudinal distance between 20 S and 44 S


1
Expert's answer
2022-01-24T17:06:04-0500

The distance is

"D=(C_2-C_1)·2\pi R_E/360,\\
D=(44-20)·2\pi·6375/360=2670\text{ km}."